OpenStudy (anonymous):

solve for x tan^2x=1/3

OpenStudy (cp9454):

x=15

OpenStudy (anonymous):

the original equation was \[3\sin ^{2}x= \cos ^{2}x, 0\le x < 2\pi\]

OpenStudy (cp9454):

its correct. x=15

OpenStudy (anonymous):

and I divided the whole thing by \[\cos ^{2}x\] to get \[3\tan ^{2}x =1\]

OpenStudy (anonymous):

I just don't understand how you are getting 15

OpenStudy (anonymous):

i am pretty sure the directions want the answer to be in radians.